# Unsupervised Constitutive Model Discovery from Sparse and Noisy Data

### v1

@@ Abstract
Recently, unsupervised constitutive model discovery has gained attention through frameworks based on the Virtual Fields Method (VFM), most prominently the EUCLID approach. However, the performance of VFM-based approaches, including EUCLID, is affected by measurement noise and data sparsity, which are unavoidable in practice. The statistical finite element method (statFEM) offers a complementary perspective by providing a Bayesian framework for assimilating noisy and sparse measurements to reconstruct the full-field displacement response, together with quantified uncertainty. While statFEM recovers displacement fields under uncertainty, it does not strictly enforce consistency with constitutive relations or aim to yield interpretable constitutive models. In this work, we couple statFEM with unsupervised constitutive model discovery in the EUCLID framework, yielding statFEM–EUCLID. The framework is demonstrated for isotropic hyperelastic materials. The results show that this integration reduces sensitivity to noise and data sparsity, while ensuring that the reconstructed fields remain consistent with both equilibrium and constitutive laws.
@@ 1 Introduction
The vast number of constitutive models for solids proposed in the literature immediately raises a central question: how can one select the most appropriate model for a given dataset? Traditionally, this question of model selection is resolved through expert knowledge and iterative trial and error, a process that can be very time-consuming. This is because constitutive modeling is accompanied by the problem of parameter identification: once a constitutive law is postulated, its parameters must be calibrated against the experimental data. The best model is then typically the model that best explains the given data. Constitutive model discovery aims to automate this process by jointly determining both the model structure and its parameters directly from data. Recent approaches range from sparsity-promoting regression methods [1, 2, 3], to symbolic regression [4], to neural-network-based formulations; see [5, 6] for comprehensive reviews.
@@
Modern full-field experimental techniques such as Digital Image Correlation (DIC) provide displacement and strain fields with high spatial resolution, even under complex loading scenarios [7, 8, 9, 10]. However, stresses are, by their nature, not directly observable and are impossible to infer, especially under heterogeneous or non-proportional loadings. This precludes the direct use of supervised model discovery, which requires paired stress–strain data [11], and motivated the development of unsupervised discovery approaches [1]. Unsupervised discovery constitutes an inverse problem, which is composed of a state equation (given by the boundary value problem) and an observation equation that measures the distance between simulations and data.
@@
Finite Element Model Updating (FEMU) [12, 13, 14] casts the identification problem as a PDE-constrained optimization task, where material parameters are iteratively updated to minimize the discrepancy between simulated and measured displacement fields. In practice, FEMU is typically implemented as a Nonlinear Least-Squares FEM approach, where the mismatch between simulation and measurement is minimized through iterative solvers. While this formulation is highly flexible and can accommodate arbitrary constitutive forms, it requires costly repeated forward finite element solves. The recently proposed Automatically Differentiable Model Updating (ADiMU) framework [13] enhances computational efficiency by utilizing automatic differentiation to compute parameter sensitivities, thereby eliminating the need for finite-difference approximations.
@@
The Virtual Fields Method (VFM) [15] provides a variational framework that does not require any forward solves. Here, the nodal degrees of freedom are equated with the observational data, and parameters are identified such that the state equation is satisfied within a norm. Building on this principle, the unsupervised constitutive model discovery framework known as EUCLID [1, 16] formulates the discovery task as a sparse regression problem on model libraries constructed from, e.g., Ogden or generalized Mooney-Rivlin features to describe hyperelastic materials, directly enforcing equilibrium in weak form. This regression-based perspective enables the automated selection of parsimonious closed-form constitutive models. Bayesian extensions of EUCLID [17] further quantify uncertainty, while neural-network-based formulations [18, 19, 20] have demonstrated the potential of unsupervised deep learning for constitutive identification. Variational System Identification (VSI) [21, 22], also inspired by VFM, offers a different weak-form strategy, where global sparse regression is replaced with stepwise elimination based on statistical hypothesis testing, thereby pruning insignificant terms from the candidate set.
@@
Despite these advances, the practical deployment of unsupervised approaches is fundamentally constrained by data quality. Noise amplification during numerical or automatic differentiation is well known to degrade parameter identification [23, 24]. In particular, the VFM and, consequently, EUCLID assume access to dense, continuous data and therefore struggle in the presence of severe sparsity or gappy data. To improve robustness, some recent discovery frameworks have introduced Bayesian sparsity priors [25] or ensemble-based updating strategies [26], which can stabilize the identification of partial differential equations (PDEs) from noisy or incomplete data. While promising in the context of general PDEs identification, these strategies do not address the mechanics-specific challenge of constitutive model discovery.
@@
A more systematic treatment of uncertainty has long been established in the uncertainty quantification community, where iterative Bayesian assimilation frameworks are used to deal with imperfect data. By treating model calibration as a sequential inference problem, such approaches refine predictions by continuously updating probabilistic beliefs in light of noisy observations [27, 28, 29]. The statistical finite element method (statFEM) [30, 31, 32, 33] is a prominent example: it reconstructs full-field displacements from sparse and noisy measurements by applying Bayesian inference directly into the finite element solution. Recent works have established polynomial convergence guarantees [34] and probabilistic displacement reconstructions under polynomial chaos expansions [35]. Yet, while statFEM robustly recovers displacement fields under uncertainty, it neither aims to strictly enforce constitutive consistency nor to provide interpretable material models.
@@
The remainder of this manuscript is organized as follows. In Sec. 2, the proposed statFEM–EUCLID framework is developed from rigorous theory, following the notion of reduced, intermediate, and all-at-once approaches for parameter identification and discovery [6]. Sec. 3 presents the governing equations of the boundary value problem, the construction of the constitutive model library, and the polynomial chaos expansion used for uncertainty propagation, i.e., the computation of forecasted displacements. The forecasted, stochastic displacement field is assimilated with sparse, noisy measurements as detailed in Sec. 4, and the subsequent discovery step is outlined in Sec. 5. Numerical examples in Sec. 6 demonstrate the effectiveness of the framework for hyperelastic model discovery from sparse and noisy data. Finally, conclusions and perspectives are provided in Sec. 7.
@@ 2 Unsupervised Constitutive Model Discovery
Let $\mathbf{u}$ denote the displacement field and $\bm{\kappa}\in\mathbb{R}^{n_{\phi}}$ a vector of material parameters that obey the state and observation equations
@@
$\mathbf{y}=\boldsymbol{\mathbf{O}}(\mathbf{u})+\mathbf{e}$, (2)
@@
The problem stated in (3) is referred to as the all-at-once approach (AAO) and is rarely used in practice. Instead, a deterministic variational identification strategy rooted in the VFM is employed [36, 15]. The VFM can be interpreted as a limiting case of (3), where the data misfit weight $\omega_{e}\to\infty$, i.e., the measured kinematic field $\mathbf{y}$ is treated as exact. Thus, the state equation is minimized directly for the material parameters, and we obtain:
[…]
In the context of constitutive model discovery, one seeks to identify a sparse parameter vector $\bm{\kappa}^{*}\in\mathbb{R}^{n_{\phi}^{\mathcal{A}}}$, with $n_{\phi}^{\mathcal{A}}<<n_{\phi}$. Sparsity can be enforced, for instance, using $\ell_{p}$-regularization, which has been promoted in the context of EUCLID [1]. In the present contribution, we consider an $\ell_{1}$- or LASSO-regularized [37] version of the VFM (4), which yields:
@@
    Uncertainty propagation: For given material parameters $\bm{\kappa}$ and uncertain boundary traction $\overline{\boldsymbol{\mathbf{t}}}_{\text{N}}(\xi)$, a stochastic boundary value problem is solved (see Sec. 3). Its solution is referred to as the forecasted displacement field $\mathbf{u}^{\text{f}}(\xi)$ that obeys the state equation:
@@
    Bayesian state update (statFEM): Given the observational data $\mathbf{y}$ and the forecast $\mathbf{u}^{\text{f}}$, the observation model (2) is used to perform a Bayesian update of the displacement field using statFEM. The assimilated state $\mathbf{u}^{\text{a}}$ with posterior density $f_{\mathbf{u}^{\text{a}}}$ determined using the observation equation (2) yielding:
@@
    Here, $\bm{\mu}_{\mathbf{u}^{\text{a}}}$ and $\boldsymbol{\mathbf{C}}_{\mathbf{u}^{\text{a}}}$ denote the posterior mean and covariance of the assimilated state, see Sec. 4 for details.
@@
    Convergence check: The Root Mean Squared Error (RMSE) between assimilated state $\bm{\mu}_{\mathbf{z}}=\boldsymbol{\mathbf{O}}(\bm{\mu}_{\mathbf{u}^{\text{a}}})$ at sensor locations and data $\mathbf{y}$ is defined as:
@@
    The assimilated displacement $\mathbf{u}^{\text{a}}$ obeys the stochastic observation equation (2). This allows us to equate the state $\mathbf{u}$ with the assimilated mean ($\mathbf{u}=\bm{\mu}_{\mathbf{u}^{\text{a}}}$) in the deterministic VFM or EUCLID formulations. After this step, we go back to Step 1 using $\bm{\kappa}^{*}=\bm{\kappa}_{\lambda^{*}}$, see Sec. 5.
If convergence cannot be reached, additional data can be introduced to improve identifiability and robustness - either by increasing the number of sensors $n_{\text{sen}}$, or by increasing the number of realizations per sensor $n_{\text{r}}$. The effect of $n_{\text{sen}}$ and $n_{\text{r}}$ has been studied, e.g., in [30, 33, 35]. In this contribution, we restrict ourselves to cases with $n_{\text{r}}=1$ and examine the effect of $n_{\text{sen}}$.
Fig. 1: Graphical illustration of the iterative model discovery framework. The loop iteratively updates the initial model predictions based on observational data via statFEM and discovers a constitutive model until convergence is reached. The dashed line shows the displacement assimilation process without updating the constitutive model. This framework is adapted from [29].
@@ 3 Forward Problem and Uncertainty Propagation
Throughout this work, we consider the balance of linear momentum subject to Dirichlet and Neumann boundary conditions as the state equation. Both the deterministic and stochastic boundary value problems are defined in Sec. 3.1. These problems are complemented by a constitutive model library introduced in Sec. 3.2. In Sec. 3.3, the PCE-based surrogate model of the stochastic boundary value problem is discussed.
@@ 3.1 Boundary Value Problem
This abstract formulation accommodates various types of constitutive behavior. Herein, we consider a linear elastic model
@@
with $\bm{\varepsilon}=\tfrac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{\top})$ the linear strain measure and $\mathbb{C}$ the isotropic elasticity tensor defined by Young’s modulus $E$ and Poisson ratio $\nu$. In addition, we consider hyperelastic models from a library introduced in Sec. 3.2. The discretization of the DBVP is discussed in the context of EUCLID in Sec. 5.
In forward simulations involving uncertainty, we consider stochasticity only in the external loading, not in the material model or its parameters. Thus, the traction $\overline{\boldsymbol{\mathbf{t}}}$ is modeled as a random field parameterized by a standard Gaussian variable $\xi$. This yields the stochastic boundary value problem (SBVP):
@@
Herein, the SBVP is approximated via PCE, see Sec. 3.3.

### v2

@@ Abstract
Recently, unsupervised constitutive model discovery has gained attention through frameworks based on the Virtual Fields Method (VFM), most prominently the EUCLID approach. However, the performance of VFM-based approaches, including EUCLID, is affected by measurement noise and data sparsity, which are unavoidable in practice. The statistical finite element method (statFEM) offers a complementary perspective by providing a Bayesian framework for assimilating noisy and sparse measurements to reconstruct the full-field displacement response, together with quantified uncertainty. While statFEM recovers displacement fields under uncertainty, it does not strictly enforce consistency with constitutive relations. In this work, we integrate statFEM with unsupervised constitutive model discovery in the EUCLID framework, yielding statFEM–EUCLID. The framework is demonstrated for isotropic hyperelastic materials. The results show that this integration reduces sensitivity to noise and data sparsity, while ensuring that the reconstructed fields remain consistent with both equilibrium and constitutive laws.
@@ 1 Introduction
The vast number of constitutive models for solids proposed in the literature immediately raises a central question: how can one select the most appropriate model for a given dataset? Traditionally, this question of model selection is resolved through expert knowledge and iterative trial and error, a process that can be very time-consuming. This is because constitutive modeling is accompanied by the problem of parameter identification: once a constitutive law is postulated, its parameters must be calibrated against the experimental data. The best model is then typically the model that best explains the given data. Constitutive model discovery aims to automate this process by jointly determining both the model structure and its parameters directly from data. Recent approaches range from sparsity-promoting regression methods [flaschel2021unsupervised, flaschel2023automatedbrain, urre2025MDisc], to symbolic regression [kissas2024language], to neural-network-based formulations; see [fuhg2024review, romer2025reduced] for comprehensive reviews.
@@
Modern full-field experimental techniques such as Digital Image Correlation (DIC) provide displacement and strain fields with high spatial resolution, even under complex loading scenarios [sutton2008DIC, pierron2023MatTesting2, he2024review, bogusz2024application]. However, stresses are, by their nature, not directly observable and are impossible to infer, especially under heterogeneous or non-proportional loadings. This precludes the direct use of supervised model discovery, which requires paired stress–strain data [tu2020stress], and motivated the development of unsupervised discovery approaches [flaschel2021unsupervised]. Unsupervised discovery constitutes an inverse problem, which is composed of a state equation (given by the boundary value problem) and an observation equation that measures the distance between simulations and data.
@@
Finite Element Model Updating (FEMU) [chen2024finite, akerson2025learning] casts the identification problem as a PDE-constrained optimization task, where material parameters are iteratively updated to minimize the discrepancy between simulated and measured displacement fields. In practice, FEMU is typically implemented as a Nonlinear Least-Squares FEM approach, where the mismatch between simulation and measurement is minimized through iterative solvers. While this formulation is highly flexible and can accommodate arbitrary constitutive forms, it requires costly repeated forward finite element solves. The most efficient implementation is via adjoints, where, independently of the number of material parameters, only one forward solve per iteration is required. When numerical differentiation via finite differences is used to approximate the derivative of the loss with respect to material parameters, the number of necessary forward calls depends on the order of the finite difference stencil and the number of material parameters. Automatic differentiation can be used to reduce the cost, see e.g. [seidl2022calibration].
@@
The Virtual Fields Method (VFM) [pierron2012virtual] provides a variational framework that does not require any forward solves. Here, the nodal degrees of freedom are equated with the observational data, and parameters are identified such that the state equation is satisfied within a norm. Building on this principle, the unsupervised constitutive model discovery framework known as EUCLID [flaschel2021unsupervised, flaschel2023automated] formulates the discovery task as a sparse regression problem on model libraries constructed from, e.g., Ogden or generalized Mooney-Rivlin features to describe hyperelastic materials, directly enforcing equilibrium in weak form. This regression-based perspective enables the automated selection of parsimonious closed-form constitutive models. Bayesian extensions of EUCLID [joshi2022bayesian] further quantify uncertainty, while neural-network-based formulations [thakolkaran2022nn, lourencco2024indirect, li2022equilibriumCNN] have demonstrated the potential of unsupervised deep learning for constitutive identification. Variational System Identification (VSI) [wang2021inference, wang2021variational], also inspired by VFM, offers a different weak-form strategy, where global sparse regression is replaced with stepwise elimination based on statistical hypothesis testing, thereby pruning insignificant terms from the candidate set. Recently, a database-driven strategy called Material Fingerprinting (MF) [flaschel2025material] was presented. MF reinterprets the discovery problem as a pattern recognition task, matching experimental data against a precomputed database of stress responses from a finite set of established constitutive models (e.g., Gent, Ogden). This avoids online optimization, but comes at the cost of constructing a comprehensive database.
@@
While a recent study [abbasi2025discovery] has demonstrated that EUCLID can be applied to experimental full-field displacement data, VFM-based approaches inherently require dense spatial data to compute deformation gradients. They cannot be applied directly to sparse point measurements (e.g., only a few sensors). The practical deployment of unsupervised approaches is further constrained by data quality. Noise amplification during numerical or automatic differentiation is well known to degrade parameter identification [avril2004sensitivity, rossi2015effect]. To improve robustness, some recent discovery frameworks have introduced Bayesian sparsity priors [hirsh2022sparsifying] or ensemble-based updating strategies [fasel2022ensemble], which can stabilize the identification of partial differential equations (PDEs) from noisy or incomplete data. While promising in the context of general PDEs identification, these strategies do not address the mechanics-specific challenge of constitutive model discovery.
@@
A more systematic treatment of uncertainty has long been established in the uncertainty quantification community, where iterative Bayesian assimilation frameworks are used to deal with imperfect data. By treating model calibration as a sequential inference problem, such approaches refine predictions by continuously updating probabilistic beliefs in light of noisy observations [kennedy2001bayesian, jiang2015surrogate, Arendt2012]. The statistical finite element method (statFEM) [girolami2021statistical, akyildiz2022statistical, koh2023stochastic, narouie2023inferring] is a prominent example: it reconstructs full-field displacements from sparse and noisy measurements by applying Bayesian inference directly into the finite element solution. Recent works have established polynomial convergence guarantees [karvonen2025error] and probabilistic displacement reconstructions under polynomial chaos expansions (PCE) [narouie2025mechanical]. Yet, while statFEM robustly recovers displacement fields under uncertainty, it neither aims to strictly enforce constitutive consistency nor to provide interpretable material models.
@@
The remainder of this manuscript is organized as follows. In Sec. 2, the proposed statFEM–EUCLID framework is developed from rigorous theory, following the notion of reduced, intermediate, and all-at-once approaches for parameter identification and discovery [romer2025reduced]. Sec. 3 presents the governing equations of the boundary value problem, the construction of the constitutive model library, and the PCE used for uncertainty propagation, i.e., the computation of forecasted displacements. The forecasted, stochastic displacement field is assimilated with sparse, noisy measurements as detailed in Sec. 4, and the subsequent discovery step is outlined in Sec. 5. Numerical examples in Sec. 6 demonstrate the effectiveness of the framework for hyperelastic model discovery from sparse and noisy data. Finally, conclusions and perspectives are provided in Sec. 7.
@@ 2 Unsupervised Constitutive Model Discovery
Let $\mathbf{u}\in\mathbb{R}^{n_{\text{gdof}}}$ denote the displacement vector and $\bm{\kappa}\in\mathbb{R}^{n_{\phi}}$ a vector of material parameters. The inverse problem is governed by the state and observation equations:
@@
$\mathbf{y}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\mathbf{H}}}\,\mathbf{u}+\mathbf{e}$. (2)
@@
In a deterministic setting, we can identify both the unknown displacements $\mathbf{u}$ and material parameters $\bm{\kappa}$ such that the $\ell_{2}$ norms of state (1) and observation (2) are minimized, c.f. [romer2025reduced]:
@@
Here, the noise is considered negligible ($\mathbf{e}=\mathbf{0})$ and $\omega_{s}$ and $\omega_{e}$ control the relative importance between physics (state equation) and data (observation equation). Note that, along the same lines, the state and observation equations (1) and (2) can be considered as stochastic. In this case, under suitable assumptions, (3) can be derived from a Maximum a Posteriori (MAP) estimate, see Appx. B for details.
The problem stated in (3) is referred to as the all-at-once approach (AAO) and is rarely used in practice. Instead, the reduced approach with $\omega_{s}\rightarrow\infty$ is most frequently employed, i.e. the state equation is treated as exact. The FEMU approach falls under this category, see e.g. [romer2025reduced] and references therein. Herein, we employ an intermediate approach with $\omega_{e}\to\infty$, where the observation equation is treated as exact. This concept is employed in the VFM [grediac2006virtual, pierron2012virtual]. Here, the state equation is minimized directly for the material parameters, and we obtain:
@@
In the context of constitutive model discovery, one seeks to identify a sparse parameter vector $\bm{\kappa}^{*}\in\mathbb{R}^{n_{\phi}^{\mathcal{A}}}$, with $n_{\phi}^{\mathcal{A}}<<n_{\phi}$. Sparsity can be enforced, for instance, using $\ell_{p}$-regularization, which has been promoted in the context of EUCLID [flaschel2021unsupervised]. In the present contribution, we consider an $\ell_{1}$- or LASSO-regularized [mcculloch2024LPRegularizationModelDiscovery] version of the VFM (4), which yields:
@@
    Uncertainty propagation: For given material parameters $\bm{\kappa}$ and uncertain boundary traction $\overline{\boldsymbol{\mathbf{t}}}_{\text{N}}(\xi)$, a stochastic boundary value problem is solved (see Sec. 3). Its solution is referred to as the nodal forecasted displacement field $\mathbf{u}^{\text{f}}\in\mathbb{R}^{n_{\text{gdof}}}$ that follows from the state equation:
@@
    Bayesian state update (statFEM): Given the observational data $\mathbf{y}$ and the forecast $\mathbf{u}^{\text{f}}$, the observation model (2) is used to perform a Bayesian update of the displacement field using statFEM. The assimilated nodal state $\mathbf{u}^{\text{a}}\in\mathbb{R}^{n_{\text{gdof}}}$, with posterior density $f_{\mathbf{u}^{\text{a}}}$, is determined from the observation equation (2), yielding:
@@
    Here, $\bm{\mu}_{\mathbf{u}^{\text{a}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathbb{R}^{n_{\text{gdof}}}}$ and $\boldsymbol{\mathbf{C}}_{\mathbf{u}^{\text{a}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathbb{R}^{n_{\text{gdof}}\times n_{\text{gdof}}}}$ denote the posterior mean and covariance of the assimilated state, see Sec. 4 for details.
@@
    Convergence check: The Root Mean Squared Error (RMSE) between assimilated state $\bm{\mu}_{\mathbf{z}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\mathbf{H}}}\,\bm{\mu}_{\mathbf{u}^{\text{a}}}$ at sensor locations and data $\mathbf{y}$ is defined as:
@@
    The assimilated mean displacement $\bm{\mu}_{\mathbf{u}^{\text{a}}}$ is considered the approximate solution of the stochastic observation equation (2). This viewpoint allows us to equate the state $\mathbf{u}$ with the assimilated mean ($\mathbf{u}=\bm{\mu}_{\mathbf{u}^{\text{a}}}$) in the deterministic VFM or EUCLID formulations. After this step, we go back to Step 1 using $\bm{\kappa}^{*}=\bm{\kappa}_{\lambda^{*}}$, see Sec. 5.
If convergence cannot be reached, additional data can be introduced to improve identifiability and robustness - either by increasing the number of sensors $n_{\text{sen}}$, or by increasing the number of realizations per sensor $n_{\text{r}}$. The effect of $n_{\text{sen}}$ and $n_{\text{r}}$ has been studied, e.g., in [girolami2021statistical, narouie2023inferring, narouie2025mechanical]. In this contribution, we restrict ourselves to cases with $n_{\text{r}}=1$ and examine the effect of $n_{\text{sen}}$.
Fig. 1: Graphical illustration of the iterative model discovery framework. The loop iteratively updates the initial model predictions based on observational data via statFEM and discovers a constitutive model until convergence is reached. The dashed line shows the displacement assimilation process without updating the constitutive model. This framework is adapted from [Arendt2012].
@@ 3 Forward Problem and Uncertainty Propagation
Throughout this work, we consider the balance of linear momentum subject to Dirichlet and Neumann boundary conditions as the state equation. Both the deterministic and stochastic boundary value problems are defined in Sec. 3.1. These problems are complemented by a constitutive model library introduced in Sec. 3.2. In Sec. 3.3, the Polynomial-Chaos-based surrogate model of the stochastic boundary value problem is discussed.
@@ 3.1 Boundary Value Problem
Here, with a slight abuse of notation, $\mathbf{u}\in\mathbb{R}^{3}$ denotes the displacement field rather than the nodal displacement vector. The abstract formulation (11) accommodates various types of constitutive behavior. Herein, we consider a linear elastic model
@@
In forward simulations involving uncertainty, we consider stochasticity only in the external loading, not in the material model or its parameters. Thus, the traction $\overline{\boldsymbol{\mathbf{t}}}(\xi)$ is parametrized with a standard Gaussian random variable $\xi$. This yields the stochastic boundary value problem (SBVP):
[…]
Notice that the DBVP (11) leads to the VFM, whereas the SBVP (13) is required in the statFEM step to compute the forecasted displacement. Herein, the solution of the SBVP is approximated via PCE, see Sec. 3.3.